OpenStudy (anonymous):

Solve for x. 3x − 5 = 4 A. x = -0.33 B. x = 8.66 C. x = 2 D. x = 3

OpenStudy (uri):

Add 5 both sides 3x − 5 = 4 +5 +5 ========= 3x =9

OpenStudy (uri):

@heybhai Tell her how to do it...o_O

OpenStudy (uri):

Also this is not the right way you are doing..Are you pakistani? @heybhai

OpenStudy (anonymous):

@uri are u racist?

OpenStudy (uri):

I used to do that in the same way but someone yesterday told me that we get the same answers but the steps are not correct and you will lose marks in test .No i ain't racist.I'm a Pakistani myself.Teachers don't teach this in the correct way here.

OpenStudy (uri):

The correct way is the one i showed.

OpenStudy (uri):

I learnt it yesterday myself.

OpenStudy (uri):

From where ever you are that's not the correct way,I wasn't rude to you..so please be nice. :)